On Pakistan Day 9 kg sweets were distributed among children. each child receives as more grams as the number of children and still 164 grams of sweets were left over. find number of children.
Answer by Theo(13342)

9 kilograms = 9000 grams.

each child receives as many grams as the number of children and there are 164 grams of sweets left over.

take the 164 grams left over and you have 9000 - 164 = 8836 grams were consumed.

let x = number of children.
let y = number of grams per child.

you have x*y = 8836

since the number of grams per child is the same as the number of children, then you have y = x and the equation of x*y = 8836 becomes x^2 = 8836.

solve for x to get x = sqrt(8836) = 96.

there were 96 children and each child received 96 grams of sweets.

96 * 96 = 8836 sweets consumed by the children plus 164 remaining.